OpenStudy (anonymous):

How do computers evaluate functions such as the sine and cosine function, or the inverse sine/inverse cosine function?

OpenStudy (anonymous):

using taylor series

OpenStudy (anonymous):

\[\sin(x) = \sum_{i=0}^{\infty}(-1)^{n}\frac{x^{n+1}}{(n+1)!}\]

OpenStudy (anonymous):

yes, like that

OpenStudy (anonymous):

so in a programming language, that would be an iterative loop. When does it stop?

OpenStudy (anonymous):

the processor stops when the solution of the previous step is the binary equivalent of the last step, or in other words you have reached the 32bit/64bit/128bit answer for your math library. You can install 128bit math libraries...